# SpikCommander: A High-performance Spiking Transformer with Multi-view Learning for Efficient Speech Command Recognition

### v2

@@ Abstract
Spiking neural networks (SNNs) offer a promising path toward energy-efficient speech command recognition (SCR) by leveraging their event-driven processing paradigm. However, existing SNN-based SCR methods often struggle to capture rich temporal dependencies and contextual information from speech due to limited temporal modeling and binary spike-based representations. To address these challenges, we first introduce the multi-view spiking temporal-aware self-attention (MSTASA) module, which combines effective spiking temporal-aware attention with a multi-view learning framework to model complementary temporal dependencies in speech commands. Building on MSTASA, we further propose SpikCommander, a fully spike-driven transformer architecture that integrates MSTASA with a spiking contextual refinement MLP (SCR-MLP) to jointly enhance temporal context modeling and channel-wise feature integration. We evaluate our method on three benchmark datasets: the Spiking Heidelberg Dataset (SHD), the Spiking Speech Commands (SSC), and the Google Speech Commands V2 (GSC). Extensive experiments demonstrate that SpikCommander consistently outperforms state-of-the-art (SOTA) SNN approaches with fewer parameters under comparable time steps, highlighting its effectiveness and efficiency for robust speech command recognition. Code and Appendix are available at https://github.com/JackieWang9811/SpikCommander.
@@ Introduction
Recognized as the third generation of neural networks (maass1997networks), spiking neural networks (SNNs) effectively mimic the spiking behavior of biological neural circuits through binary information communication (guo2024spgesture). This brain-inspired attribute enables SNNs to achieve high computational efficiency with their event-driven processing paradigm and significantly reduce energy consumption through spike-based accumulation (AC) operations. These advantages become especially prominent when SNNs are deployed on neuromorphic hardware platforms such as Tianjic (pei2019towards), Loihi (davies2021advancing), and TrueNorth (akopyan2015truenorth), positioning them as a compelling and energy-efficient alternative to conventional artificial neural networks (ANNs) (fang2023spikingjelly).
Recent studies (wang2025efficient; shen2024tim; hammouamrilearning2024; wang2024global; he2024msat) have demonstrated that SNN models can effectively tackle speech command recognition (SCR) tasks (also known as keyword spotting) by leveraging temporal-embedded information through their spatio-temporal encoding mechanisms. These characteristics make SNNs promising candidates for crucial auditory front-end applications, where the goal is to efficiently detect predefined spoken commands in dynamic, real-world settings. Despite these merits, current SNN-based approaches for speech command recognition still tend to underperform compared to ANNs (gong2021ast; schone2024scalable), largely due to the challenges posed by their binary and sparse spike-based representations, which hinder the effectiveness of conventional operations over continuous-valued features for capturing complex speech patterns.
@@
To address the aforementioned challenges of applying SNNs to speech command recognition, we structure our work around two key parts. First, although recent spiking attention mechanisms have achieved notable success in vision (zhou2023spikformer; yao2023spikedriven; hua2025msvit) and language modeling (xing2024spikellm; xing2024spikelm; zhang2024spikingminilm), they remain underexplored for SCR tasks. To bridge this gap, we introduce a spiking temporal-aware self-attention (STASA), an efficient module with linear complexity designed to model essential temporal dependencies in speech sequences. We further extend it into a multi-view learning framework (MSTASA) that captures complementary temporal cues via three paths: a sliding-window STASA branch for local context, a long-range STASA branch for global context, and a convolutional path (termed V-branch) that operates on the value stream to inject shift-invariant, position-aware patterns. This unified design facilitates comprehensive temporal modeling. Second, based on MSTASA, we present SpikCommander, a compact spiking transformer architecture that integrates MSTASA with a spiking contextual refinement MLP (SCR-MLP). SCR-MLP adopts a channel-mixing MLP structure with selective contextual refinement to enhance both context modeling and inter-channel interactions. Together, SpikCommander produces expressive spike-based representations for SCR tasks.
We evaluate our SpikCommander on three speech command datasets: the Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC), both of which are spiking command datasets (cramer2020heidelberg), as well as the Google Speech Commands V2 (GSC) (warden2018speech). Our experimental results demonstrate that SpikCommander outperforms current SOTA SNN methods across these datasets with fewer parameters under the same time step settings.
@@
Recent efforts have sought to extend Transformer architectures (vaswani2017attention; dosovitskiy2020image) to the SNNs. As illustrated in Fig.1, existing approaches broadly fall into two categories based on their architecture. Fig.1(a) illustrates models that adopt spiking self-attention mechanisms for global context modeling, combined with channel-wise MLPs for feature mixing. Such as Spikformer (zhou2023spikformer; zhou2023spikingformer), which introduces spiking self-attention (SSA), and Spike-driven Transformer (SDT) (yao2023spikedriven; yao2024spikedriven), with a variant called spike-driven self-attention (SDSA). Fig.1(b) shows hybrid designs integrating spiking attention with convolutional operations to separate global and local processing pathways, as shown in SpikeSCR (wang2025efficient) and SGLFormer (zhang2024sglformer). In contrast, our proposed SpikCommander (see Fig. 1(c)) introduces two key innovations: i) a unified multi-view learning module (MSTASA) that leverages sliding-window STASA, long-range STASA, and a V-branch to jointly model global and local dependencies; ii) a spiking contextual refinement MLP (SCR-MLP) that explicitly refines contextual and channel-wise features. These innovations enable comprehensive feature modeling, clearly distinguishing SpikCommander from prior designs.
@@
Recent research can be broadly grouped into four directions. First, innovative spiking neuron designs have been proposed to improve the learning of temporal dynamics, such as d-cAdLIF (deckers2024co) and SE-adLIF (baronig2025advancing). The second focuses on developing delay-learning based methods that introduce learnable delay mechanisms to enrich temporal representations, as seen in DL-SNN (sun2023learnable) and DCLS-Delays (hammouamrilearning2024). Third, spiking-based memory modules have been employed to better capture sequential dependencies, including DH-SNN (zheng2024temporal) and Spiking LMUFormer (liu2024lmuformer). Finally, recent work involves incorporating spiking attention mechanisms for SCR, such as TIM (shen2024tim), SpikeSCR (wang2025efficient), and PfA-SNN (sun2025towards). Despite recent progress, the inherent binary and sparse nature of spikes still limits effective feature extraction from speech commands. Our SpikCommander addresses this with a more expressive and energy-efficient architecture.
@@ Methods
Spiking neurons are a fundamental component of SNNs, offering bio-plausible abstractions of neuronal dynamics (ma2025spiking). We adopt the Leaky Integrate-and-Fire (LIF) neuron model, which has been widely validated for its effectiveness in spiking architectures (roy2019towards). The dynamics of the LIF neuron can be described as:
@@
STASA readily extends to multi-head attention by splitting features dimension ($D$) and applying temporal-aware summation per head. Details of the sliding-window variant of STASA are provided in Appendix C. Compared with the classic spiking self-attention (SSA) (zhou2023spikformer) that relies on the matrix multiplication $Q_{S}K_{S}^{\top}$ with $\mathcal{O}(N^{2}D)$ complexity, where $N$ denotes the number of tokens, which corresponds to the number of time steps $T$ in SCR tasks. STASA reduces the computational cost to linear complexity $\mathcal{O}(ND)$, making it more efficient for long time sequences.
@@
SHD SDT (1L) (yao2023spikedriven) 1.77 100 89.61
Spikformer (1L) (zhou2023spikformer) 1.77 100 90.10
DH-SNN (2L) (zheng2024temporal) 0.05 1000 92.10
d-cAdLIF (2L) (deckers2024co) 0.08 100 94.85
DCLS-Delays (2L) (hammouamrilearning2024) 0.20 100 95.07
SpikeSCR (1L) (wang2025efficient) 0.26 100 95.60
SE-adLIF (2L) (baronig2025advancing) 0.45 250 95.81
@@
SSC SDT (2L) (yao2023spikedriven) 2.57 100 79.82
DCLS-Delays (2L) (hammouamrilearning2024) 1.40 100 80.16
Spikformer (2L) (zhou2023spikformer) 2.57 100 80.18
Pfa-SNN (sun2025towards) 0.71 100 80.18
d-cAdLIF (2L) (deckers2024co) 0.70 100 80.23
SE-adLIF (2L) (baronig2025advancing) 1.60 250 80.44
DCLS-Delays (3L) (hammouamrilearning2024) 2.50 100 80.69
DH-SNN (3L) (zheng2024temporal) 0.35 1000 82.46
SpikeSCR (1L) (wang2025efficient) 1.71 100 82.54
SpikeSCR (2L) (wang2025efficient) 3.30 100 82.79
@@
GSC Spikformer (2L) (zhou2023spikformer) 2.57 100 91.86
SDT (2L) (yao2023spikedriven) 2.57 100 91.88
DH-SNN (3L) (zheng2024temporal) 0.11 1000 94.05
DCLS-Delays (2L) (hammouamrilearning2024) 1.40 100 95.00
DCLS-Delays (3L) (hammouamrilearning2024) 2.50 100 95.35
SpikeSCR (1L) (wang2025efficient) 1.71 100 95.46
SpikeSCR (2L) (wang2025efficient) 3.30 100 95.60
d-cAdLIF (2L) (deckers2024co) 0.61 100 95.69
Spiking LMUFormer (liu2024lmuformer) 1.69 96.12
@@
Note that our model is trained end-to-end from scratch using backpropagation-through-time (BPTT) with surrogate gradients (fang2021deep). Table 7 in Appendix B provides detailed settings of the optimizer and learning rate. The classification head encodes per-time-step outputs using a softmax over the spikes $s_{i}[t]$ at each time step $t$:
@@
We evaluate our models on two spiking datasets, SHD and SSC (cramer2020heidelberg), as well as the non-spiking counterpart of SSC, GSC (warden2018speech). SHD contains 10k recordings across 20 classes (digits zero to nine in English and German). SSC and GSC are larger datasets, each comprising 100k recordings with 35 distinct speech command classes. Table 6 in Appendix B summarizes detailed dataset statistics. We follow the same preprocessing pipeline as in (wang2025efficient; hammouamrilearning2024) for all three datasets. For SHD and SSC, the original 700 input neurons are first reduced to 140 by applying spatio-temporal binning over every five neurons. To standardize input lengths, all samples are zero-padded to a fixed number of time steps. Specifically, the total number of time steps $T$ is computed as $T=1000/\Delta t$, where 1000 ms is the total duration of each spiking sample, and different temporal resolutions are discretized using fixed-duration windows of $\Delta t$ ms. Considering $\Delta t\in\{1,4,10\}$ yields $T\in\{1000,250,100\}$, respectively. For GSC, all audio waveforms are first downsampled from 16 kHz to 8 kHz. We then compute Mel spectrograms with 140 frequency bins, matching the input dimensionality used for SSC. To approximately match the temporal resolutions used in SSC, we use a fixed window size $l=256$ (corresponding to 32 ms) and vary the hop length $h$ to control the number of time steps. The time steps $T$ are calculated as T=(8000l)/h+1T=(8000-l)/h+1, yielding $T\in\{1000,250,100\}$ for $h\in\{8,32,80\}$. For consistency with prior work, we focus on 100 time steps, with the sliding window radius $w$ set to 20.
@@
SpikeSCR (1L-16-256) (wang2025efficient) 1.63 8 95.56 0.011 0.019 0.067
SpikeSCR (2L-16-256) (wang2025efficient) 3.15 8 95.60 0.011 0.049 0.094
Spiking LMUFormer (liu2024lmuformer) 1.69 16 96.12 0.007 0.031 0.059
@@
To enhance robustness and generalization, we apply augmentation tailored to speech commands, focusing on two modalities: Mel spectrograms and spike trains. For Mel spectrograms, we adopt SpecAugment (park2019specaugment), applying frequency and time masking to improve robustness to variations in time-frequency patterns. For spike trains, we utilize an augmentation strategy (wang2025efficient) with two operations: drop-by-time and drop-by-neuron, randomly removing events along the temporal and neuronal dimensions to simulate realistic noise. The effects of these augmentations on the GSC and SSC datasets are illustrated in Fig. 5 and Fig. 6, respectively, with detailed implementation parameters provided in Appendix A.
@@
Comprehensive performance. We compare our method with recent SNN models across three datasets in terms of accuracy and model size. As shown in Table 1, SpikCommander consistently outperforms prior methods and establishes a new SOTA performance. Notably, it achieves superior performance using only standard LIF neurons, surpassing models employing more sophisticated neuron dynamics like SE-adLIF (baronig2025advancing) and d-cAdLIF (deckers2024co) under the same time steps. For example, on SHD, SpikCommander reaches 96.41% accuracy with just 0.19M parameters, surpassing SE-adLIF’s 95.81% with 0.45M. Moreover, delay learning-based models like DCLS-Delays (hammouamrilearning2024) have recently shown strong temporal modeling capability, yet SpikCommander achieves higher accuracy with fewer parameters at the same time steps. On SSC, it outperforms the 3-block DCLS-Delays (2.50M) by 2.57% using only 1.12M. Furthermore, when compared to spiking memory modules like DH-SNN (zheng2024temporal), SpikCommander exhibits clear advantages. Although DH-SNN achieves decent accuracy with a minimal parameter count, it relies on significantly longer sequences (e.g., 1000 time steps). SpikCommander achieves higher accuracy with only 100 time steps (e.g., 96.71% vs. 94.05% on GSC), making it more efficient for latency-sensitive scenarios. Finally, while Spikformer (zhou2023spikformer) and SDT (yao2023spikedriven) struggle with effective temporal modeling, recent SpikeSCR (wang2025efficient) achieves competitive baselines. Notably, SpikCommander achieves higher accuracy with fewer parameters. On SSC, it improves accuracy by 0.70% (83.49% vs. 82.79%) over SpikeSCR while reducing parameters by 1.18M (2.13M vs. 3.30M); and on GSC, it achieves 1.32% (96.92% vs. 95.60%) higher accuracy with the same parameter reduction. The accuracy drop of replacing STASA with SSA or SDSA validates the effectiveness of STASA and the multi-view design (see results in Table 8 in Appendix D).
@@
Efficiency on non-spiking SCR task. As shown in Table 2, we compare SpikCommander with a range of SOTA ANN and SNN models on the GSC dataset, in terms of accuracy, model size, computational operations (FLOPs/SOPs), and theoretical energy consumption. Details of the theoretical energy estimation are provided in Appendix E. Compared to ANN-based models such as AST (gong2021ast) (86.93M) and KWT-3 (berg2021keyword) (5.36M), SpikCommander achieves comparable accuracy (96.92%, under 100 time steps) while reducing energy consumption by orders of magnitude (e.g., 0.042 mJ vs. 4.84 mJ for KWT-3, 0.207 mJ for KW-MLP (morshed2021attention)). Notably, our model maintains competitive performance even at a lower sampling rate (8 kHz vs. 16 kHz), leading to reduced signal processing cost and improved efficiency. Among SNN baselines, SpikCommander also demonstrates clear advantages. Compared to Spiking LMUFormer, which attains 96.12% accuracy with 0.059 mJ energy consumption, our 2-block SpikCommander achieves +0.80% higher accuracy while consuming 28.8% less energy. Compared to SpikeSCR (2L), which achieves 95.60% accuracy with 0.094 mJ energy consumption, our 2-block SpikCommander achieves higher accuracy (+1.32%) with less than half the energy consumption (0.042 mJ).
@@
Efficiency on spiking SCR tasks. We further report a quantitative comparison of model size, synaptic operation counts (SOPs), and theoretical energy consumption on the SSC and SHD datasets under 100 time steps, as shown in Table 3. While Spikformer and SDT present moderate parameters, both models are re-implemented following the 2D spatial computation setup described in (shen2024tim), where 1D sequences are reshaped into pseudo-images and processed with spatiotemporal operations via nearest-neighbor interpolation. This leads to substantial redundant computation, resulting in significantly higher SOPs (e.g., 0.3084G on SSC). In contrast, both SpikeSCR and SpikCommander operate directly on sequential data. SpikeSCR adopts a deeper hybrid design to capture local and global dependencies, which increases its computational overhead (e.g., 0.0348G SOPs and 3.30M parameters on SSC). Notably, SpikCommander consistently outperforms SpikeSCR, achieving 42.7% lower energy consumption (0.0180 mJ vs. 0.0314 mJ), 50.3% fewer SOPs (0.0173G vs. 0.0348G), and a 35.5% smaller model size (2.13M vs. 3.30M). On SHD, SpikCommander again demonstrates superior efficiency, reducing energy consumption by 23.5%, SOPs by 39.3%, and model size by 26.9% compared to SpikeSCR. These collectively demonstrate SpikCommander’s superior trade-off between performance, parameter efficiency, and low energy consumption, underscoring its potential for low-power speech processing on neuromorphic hardware.

### v3

@@ Abstract
Spiking neural networks (SNNs) offer a promising path toward energy-efficient speech command recognition (SCR) by leveraging their event-driven processing paradigm. However, existing SNN-based SCR methods often struggle to capture rich temporal dependencies and contextual information from speech due to limited temporal modeling and binary spike-based representations. To address these challenges, we first introduce the multi-view spiking temporal-aware self-attention (MSTASA) module, which combines effective spiking temporal-aware attention with a multi-view learning framework to model complementary temporal dependencies in speech commands. Building on MSTASA, we further propose SpikCommander, a fully spike-driven transformer architecture that integrates MSTASA with a spiking contextual refinement MLP (SCR-MLP) to jointly enhance temporal context modeling and channel-wise feature integration. We evaluate our method on three benchmark datasets: the Spiking Heidelberg Dataset (SHD), the Spiking Speech Commands (SSC), and the Google Speech Commands V2 (GSC). Extensive experiments demonstrate that SpikCommander consistently outperforms state-of-the-art (SOTA) SNN approaches with fewer parameters under comparable time steps, highlighting its effectiveness and efficiency for robust speech command recognition.
Codehttps://github.com/JackieWang9811/SCommander
@@ Introduction
Recognized as the third generation of neural networks (Maass 1997), spiking neural networks (SNNs) effectively mimic the spiking behavior of biological neural circuits through binary information communication (Guo et al. 2024). This brain-inspired attribute enables SNNs to achieve high computational efficiency with their event-driven processing paradigm and significantly reduce energy consumption through spike-based accumulation (AC) operations. These advantages become especially prominent when SNNs are deployed on neuromorphic hardware platforms such as Tianjic (Pei et al. 2019), Loihi (Davies et al. 2021), and TrueNorth (Akopyan et al. 2015), positioning them as a compelling and energy-efficient alternative to conventional artificial neural networks (ANNs) (Fang et al. 2023).
Recent studies (Wang et al. 2025; Shen et al. 2024; Hammouamri et al. 2024; Wang et al. 2024b; He et al. 2024) have demonstrated that SNN models can effectively tackle speech command recognition (SCR) tasks (also known as keyword spotting) by leveraging temporal-embedded information through their spatio-temporal encoding mechanisms. These characteristics make SNNs promising candidates for crucial auditory front-end applications, where the goal is to efficiently detect predefined spoken commands in dynamic, real-world settings. Despite these merits, current SNN-based approaches for speech command recognition still tend to underperform compared to ANNs (Gong et al. 2021; Schöne et al. 2024), largely due to the challenges posed by their binary and sparse spike-based representations, which hinder the effectiveness of conventional operations over continuous-valued features for capturing complex speech patterns.
@@
To address the aforementioned challenges of applying SNNs to speech command recognition, we structure our work around two key parts. First, although recent spiking attention mechanisms have achieved notable success in vision (Zhou et al. 2023b; Yao et al. 2023; Hua et al. 2025) and language modeling (Xing et al. 2024a, b; Zhang et al. 2024c), they remain underexplored for SCR tasks. To bridge this gap, we introduce a spiking temporal-aware self-attention (STASA), an efficient module with linear complexity designed to model essential temporal dependencies in speech sequences. We further extend it into a multi-view learning framework (MSTASA) that captures complementary temporal cues via three paths: a sliding-window STASA branch for local context, a long-range STASA branch for global context, and a convolutional path (termed V-branch) that operates on the value stream to inject shift-invariant, position-aware patterns. This unified design facilitates comprehensive temporal modeling. Second, based on MSTASA, we present SpikCommander, a compact spiking transformer architecture that integrates MSTASA with a spiking contextual refinement MLP (SCR-MLP). SCR-MLP adopts a channel-mixing MLP structure with selective contextual refinement to enhance both context modeling and inter-channel interactions. Together, SpikCommander produces expressive spike-based representations for SCR tasks.
We evaluate our SpikCommander on three speech command datasets: the Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC), both of which are spiking command datasets (Cramer et al. 2020), as well as the Google Speech Commands V2 (GSC) (Warden 2018). Our experimental results demonstrate that SpikCommander outperforms current SOTA SNN methods across these datasets with fewer parameters under the same time step settings.
@@
Recent efforts have sought to extend Transformer architectures (Vaswani et al. 2017; Dosovitskiy et al. 2020) to the SNNs. As illustrated in Fig.1, existing approaches broadly fall into two categories based on their architecture. Fig.1(a) illustrates models that adopt spiking self-attention mechanisms for global context modeling, combined with channel-wise MLPs for feature mixing. Such as Spikformer (Zhou et al. 2023b, a), which introduces spiking self-attention (SSA), and Spike-driven Transformer (SDT) (Yao et al. 2023, 2024), with a variant called spike-driven self-attention (SDSA). Fig.1(b) shows hybrid designs integrating spiking attention with convolutional operations to separate global and local processing pathways, as shown in SpikeSCR (Wang et al. 2025) and SGLFormer (Zhang et al. 2024b). In contrast, our proposed SpikCommander (see Fig. 1(c)) introduces two key innovations: i) a unified multi-view learning module (MSTASA) that leverages sliding-window STASA, long-range STASA, and a V-branch to jointly model global and local dependencies; ii) a spiking contextual refinement MLP (SCR-MLP) that explicitly refines contextual and channel-wise features. These innovations enable comprehensive feature modeling, clearly distinguishing SpikCommander from prior designs.
@@
Recent research can be broadly grouped into four directions. First, innovative spiking neuron designs have been proposed to improve the learning of temporal dynamics, such as d-cAdLIF (Deckers et al. 2024) and SE-adLIF (Baronig et al. 2025). The second focuses on developing delay-learning based methods that introduce learnable delay mechanisms to enrich temporal representations, as seen in DL-SNN (Sun et al. 2023) and DCLS-Delays (Hammouamri et al. 2024). Third, spiking-based memory modules have been employed to better capture sequential dependencies, including DH-SNN (Zheng et al. 2024) and Spiking LMUFormer (Liu et al. 2024). Finally, recent work involves incorporating spiking attention mechanisms for SCR, such as TIM (Shen et al. 2024), SpikeSCR (Wang et al. 2025), and PfA-SNN (Sun et al. 2025). Despite recent progress, the inherent binary and sparse nature of spikes still limits effective feature extraction from speech commands. Our SpikCommander addresses this with a more expressive and energy-efficient architecture.
@@ Methods
Spiking neurons are a fundamental component of SNNs, offering bio-plausible abstractions of neuronal dynamics (Ma et al. 2025). We adopt the Leaky Integrate-and-Fire (LIF) neuron model, which has been widely validated for its effectiveness in spiking architectures (Roy et al. 2019). The dynamics of the LIF neuron can be described as:
@@
STASA readily extends to multi-head attention by splitting features dimension ($D$) and applying temporal-aware summation per head. Details of the sliding-window variant of STASA are provided in Appendix C. Compared with the classic spiking self-attention (SSA) (Zhou et al. 2023b) that relies on the matrix multiplication $Q_{S}K_{S}^{\top}$ with $\mathcal{O}(N^{2}D)$ complexity, where $N$ denotes the number of tokens, which corresponds to the number of time steps $T$ in SCR tasks. STASA reduces the computational cost to linear complexity $\mathcal{O}(ND)$, making it more efficient for long time sequences.
@@
SHD SDT (1L) (Yao et al. 2023) 1.77 100 89.61
Spikformer (1L) (Zhou et al. 2023b) 1.77 100 90.10
DH-SNN (2L) (Zheng et al. 2024) 0.05 1000 92.10
d-cAdLIF (2L) (Deckers et al. 2024) 0.08 100 94.85
DCLS-Delays (2L) (Hammouamri et al. 2024) 0.20 100 95.07
SpikeSCR (1L) (Wang et al. 2025) 0.26 100 95.60
SE-adLIF (2L) (Baronig et al. 2025) 0.45 250 95.81
@@
SSC SDT (2L) (Yao et al. 2023) 2.57 100 79.82
DCLS-Delays (2L) (Hammouamri et al. 2024) 1.40 100 80.16
Spikformer (2L) (Zhou et al. 2023b) 2.57 100 80.18
Pfa-SNN (Sun et al. 2025) 0.71 100 80.18
d-cAdLIF (2L) (Deckers et al. 2024) 0.70 100 80.23
SE-adLIF (2L) (Baronig et al. 2025) 1.60 250 80.44
DCLS-Delays (3L) (Hammouamri et al. 2024) 2.50 100 80.69
DH-SNN (3L) (Zheng et al. 2024) 0.35 1000 82.46
SpikeSCR (1L) (Wang et al. 2025) 1.71 100 82.54
SpikeSCR (2L) (Wang et al. 2025) 3.30 100 82.79
@@
GSC Spikformer (2L) (Zhou et al. 2023b) 2.57 100 91.86
SDT (2L) (Yao et al. 2023) 2.57 100 91.88
DH-SNN (3L) (Zheng et al. 2024) 0.11 1000 94.05
DCLS-Delays (2L) (Hammouamri et al. 2024) 1.40 100 95.00
DCLS-Delays (3L) (Hammouamri et al. 2024) 2.50 100 95.35
SpikeSCR (1L) (Wang et al. 2025) 1.71 100 95.46
SpikeSCR (2L) (Wang et al. 2025) 3.30 100 95.60
d-cAdLIF (2L) (Deckers et al. 2024) 0.61 100 95.69
Spiking LMUFormer (Liu et al. 2024) 1.69 96.12
@@
Note that our model is trained end-to-end from scratch using backpropagation-through-time (BPTT) with surrogate gradients (Fang et al. 2021). Table 7 in Appendix B provides detailed settings of the optimizer and learning rate. The classification head encodes per-time-step outputs using a softmax over the spikes $s_{i}[t]$ at each time step $t$:
@@
We evaluate our models on two spiking datasets, SHD and SSC (Cramer et al. 2020), as well as the non-spiking counterpart of SSC, GSC (Warden 2018). SHD contains 10k recordings across 20 classes (digits zero to nine in English and German). SSC and GSC are larger datasets, each comprising 100k recordings with 35 distinct speech command classes. Table 6 in Appendix B summarizes detailed dataset statistics. We follow the same preprocessing pipeline as in (Wang et al. 2025; Hammouamri et al. 2024) for all three datasets. For SHD and SSC, the original 700 input neurons are first reduced to 140 by applying spatio-temporal binning over every five neurons. To standardize input lengths, all samples are zero-padded to a fixed number of time steps. Specifically, the total number of time steps $T$ is computed as $T=1000/\Delta t$, where 1000 ms is the total duration of each spiking sample, and different temporal resolutions are discretized using fixed-duration windows of $\Delta t$ ms. Considering $\Delta t\in\{1,4,10\}$ yields $T\in\{1000,250,100\}$, respectively. For GSC, all audio waveforms are first downsampled from 16 kHz to 8 kHz. We then compute Mel spectrograms with 140 frequency bins, matching the input dimensionality used for SSC. To approximately match the temporal resolutions used in SSC, we use a fixed window size $l=256$ (corresponding to 32 ms) and vary the hop length $h$ to control the number of time steps. The time steps $T$ are calculated as T=(8000l)/h+1T=(8000-l)/h+1, yielding $T\in\{1000,250,100\}$ for $h\in\{8,32,80\}$. For consistency with prior work, we focus on 100 time steps, with the sliding window radius $w$ set to 20.
@@
SpikeSCR (1L-16-256) (Wang et al. 2025) 1.63 8 95.56 0.011 0.019 0.067
SpikeSCR (2L-16-256) (Wang et al. 2025) 3.15 8 95.60 0.011 0.049 0.094
Spiking LMUFormer (Liu et al. 2024) 1.69 16 96.12 0.007 0.031 0.059
@@
To enhance robustness and generalization, we apply augmentation tailored to speech commands, focusing on two modalities: Mel spectrograms and spike trains. For Mel spectrograms, we adopt SpecAugment (Park et al. 2019), applying frequency and time masking to improve robustness to variations in time-frequency patterns. For spike trains, we utilize an augmentation strategy (Wang et al. 2025) with two operations: drop-by-time and drop-by-neuron, randomly removing events along the temporal and neuronal dimensions to simulate realistic noise. The effects of these augmentations on the GSC and SSC datasets are illustrated in Fig. 5 and Fig. 6, respectively, with detailed implementation parameters provided in Appendix A.
@@
Comprehensive performance. We compare our method with recent SNN models across three datasets in terms of accuracy and model size. As shown in Table 1, SpikCommander consistently outperforms prior methods and establishes a new SOTA performance. Notably, it achieves superior performance using only standard LIF neurons, surpassing models employing more sophisticated neuron dynamics like SE-adLIF (Baronig et al. 2025) and d-cAdLIF (Deckers et al. 2024) under the same time steps. For example, on SHD, SpikCommander reaches 96.41% accuracy with just 0.19M parameters, surpassing SE-adLIF’s 95.81% with 0.45M. Moreover, delay learning-based models like DCLS-Delays (Hammouamri et al. 2024) have recently shown strong temporal modeling capability, yet SpikCommander achieves higher accuracy with fewer parameters at the same time steps. On SSC, it outperforms the 3-block DCLS-Delays (2.50M) by 2.57% using only 1.12M. Furthermore, when compared to spiking memory modules like DH-SNN (Zheng et al. 2024), SpikCommander exhibits clear advantages. Although DH-SNN achieves decent accuracy with a minimal parameter count, it relies on significantly longer sequences (e.g., 1000 time steps). SpikCommander achieves higher accuracy with only 100 time steps (e.g., 96.71% vs. 94.05% on GSC), making it more efficient for latency-sensitive scenarios. Finally, while Spikformer (Zhou et al. 2023b) and SDT (Yao et al. 2023) struggle with effective temporal modeling, recent SpikeSCR (Wang et al. 2025) achieves competitive baselines. Notably, SpikCommander achieves higher accuracy with fewer parameters. On SSC, it improves accuracy by 0.70% (83.49% vs. 82.79%) over SpikeSCR while reducing parameters by 1.18M (2.13M vs. 3.30M); and on GSC, it achieves 1.32% (96.92% vs. 95.60%) higher accuracy with the same parameter reduction. The accuracy drop of replacing STASA with SSA or SDSA validates the effectiveness of STASA and the multi-view design (see results in Table 8 in Appendix D).
@@
Efficiency on non-spiking SCR task. As shown in Table 2, we compare SpikCommander with a range of SOTA ANN and SNN models on the GSC dataset, in terms of accuracy, model size, computational operations (FLOPs/SOPs), and theoretical energy consumption. Details of the theoretical energy estimation are provided in Appendix E. Compared to ANN-based models such as AST (Gong et al. 2021) (86.93M) and KWT-3 (Berg et al. 2021) (5.36M), SpikCommander achieves comparable accuracy (96.92%, under 100 time steps) while reducing energy consumption by orders of magnitude (e.g., 0.042 mJ vs. 4.84 mJ for KWT-3, 0.207 mJ for KW-MLP (Morshed and Ahsan 2021)). Notably, our model maintains competitive performance even at a lower sampling rate (8 kHz vs. 16 kHz), leading to reduced signal processing cost and improved efficiency. Among SNN baselines, SpikCommander also demonstrates clear advantages. Compared to Spiking LMUFormer, which attains 96.12% accuracy with 0.059 mJ energy consumption, our 2-block SpikCommander achieves +0.80% higher accuracy while consuming 28.8% less energy. Compared to SpikeSCR (2L), which achieves 95.60% accuracy with 0.094 mJ energy consumption, our 2-block SpikCommander achieves higher accuracy (+1.32%) with less than half the energy consumption (0.042 mJ).
@@
Efficiency on spiking SCR tasks. We further report a quantitative comparison of model size, synaptic operation counts (SOPs), and theoretical energy consumption on the SSC and SHD datasets under 100 time steps, as shown in Table 3. While Spikformer and SDT present moderate parameters, both models are re-implemented following the 2D spatial computation setup described in (Shen et al. 2024), where 1D sequences are reshaped into pseudo-images and processed with spatiotemporal operations via nearest-neighbor interpolation. This leads to substantial redundant computation, resulting in significantly higher SOPs (e.g., 0.3084G on SSC). In contrast, both SpikeSCR and SpikCommander operate directly on sequential data. SpikeSCR adopts a deeper hybrid design to capture local and global dependencies, which increases its computational overhead (e.g., 0.0348G SOPs and 3.30M parameters on SSC). Notably, SpikCommander consistently outperforms SpikeSCR, achieving 42.7% lower energy consumption (0.0180 mJ vs. 0.0314 mJ), 50.3% fewer SOPs (0.0173G vs. 0.0348G), and a 35.5% smaller model size (2.13M vs. 3.30M). On SHD, SpikCommander again demonstrates superior efficiency, reducing energy consumption by 23.5%, SOPs by 39.3%, and model size by 26.9% compared to SpikeSCR. These collectively demonstrate SpikCommander’s superior trade-off between performance, parameter efficiency, and low energy consumption, underscoring its potential for low-power speech processing on neuromorphic hardware.